Circular No. 8994

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2008 T4 (HILL) R. E. Hill reports his discovery of a comet on CCD images taken with

the 0.68-m Schmidt telescope in the course of the Catalina Sky Survey (discovery observation tabulated below), a stack of four 30-s exposures showing a diffuse coma 10′′ across and a well-condensed nuclear condensation with a broad tail 20′′ long in p.a. ∼ 210^{o}. Images obtained by A. Boattini with the Mt. Lemmon 1.5-m reflector on Oct. 8.44–8.45 UT show a well-condensed coma with diameter 10′′ and a broad tail ∼ 30′′ long in p.a. 245^{o}.

2008 UT  α_{2000}  δ_{2000}  Mag.
Oct. 8.38711  1^{h}52^{m}52.^{s}10  +7^{o}39′25′′.5  17.9

The available astrometry (including prediscovery observations from Sept. 28–30), the following preliminary elliptical orbital elements, and an ephemeris appear on MPEC 2008-T90.

T = 2008 Sept.21.832 TT  ω = 331.^{o}404
e = 0.44203  Ω = 46.835 }2000.0
q = 2.44897 AU  i = 5.540

a = 4.38908 AU n^{o} = 0.107188 P = 9.20 years

2008 TC_{3}

J. Borovicka, Astronomical Institute, Czech Academy of Sciences, reports that Z. Charvat (Czech Hydrometeorological Institute) has noticed a bright spot that likely corresponds to the atmospheric entry of 2008 TC_{3} (cf. IAUC 8990) over northern Sudan on images taken by the weather satellite Meteosat 8 around Oct. 7^{d}02^{h}45^{m} UT; the spot is apparent in all twelve satellite spectral channels, spanning wavelengths 0.5–14 μm. Since the satellite takes images in scanning mode, it takes ∼ 5 min to obtain one image; consequently, the exact time of the spot’s appearance cannot be inferred easily from the image. The spot is, however, not present in the images taken at nominal times 2^{h}40^{m} and 2^{h}50^{m} UT. The geographical coordinates of the spot in the visual and near-infrared channels are λ = 32^{o}.16 east, β = +20^{o}.97 (assuming that the source of light is at sea level). The HRV channel shows an apparent tail ∼ 3 km long toward the westnorthwest. The infrared channels (displaced in the instrument focal plane and likely scanned the region 1.8 s later) show the spot at λ = 32^{o}.37 east, β = +20^{o}.89. The actual coordinates may be slightly southwest of these positions after correcting for the fireball’s altitude (∼ 30 and 22 km).

2008 October 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green